OpenStudy (anonymous):

Can someone please help me finish working this out to solve it, I have most of it but didn't finish solving it. Part 1: Explain how you calculate the restricted value(s) of x for the equation

OpenStudy (anonymous):

Here is the problem.

OpenStudy (anonymous):

and what those value(s) represent. Part 2: Explain which method (cross-multiplication or LCD) you would use to solve the equation and why. Part 3: Using complete sentences, explain the steps to solve the equation using the method you chose in step 2. Provide the solution to the equation.

OpenStudy (anonymous):

Here is what I have for an answer so far. But I only got partial credit. Part 1 - When you solve an equation for x, you have to find the values for x. The values don't follow the rules so you can't divide by zero. So you would multiply both of the sides by x 2. Part 2 - You could use either the LCD or cross multiplication method but I would use the cross multiplication method to solve the equation because I think it would be easier to use. Part 3 - (2 x)(3/2 x 1/2x)=(2 x)(4/2 x) Distribute it over, and: 3 2 x/2x = 4

OpenStudy (mathstudent55):

Cross multiplication is simply a shortcut of LCD, but it only works in one type of case: Cross multiplication only works if you have a fraction equal to a fraction. For example, if you have 20/x = 30/(x + 5) Then you cross multiply: 20(x + 5) = 30x 20x + 100 = 30x -10x = -100 x = 10 In your case, there are three fractions, so you must use LCD. To use LCD, you find the LCD of all the denominators and you multiply all fractions by the LCD. This will eliminate all denominators.

OpenStudy (anonymous):

Ok so is what I have for part #3 correct? Except for solving it? (2 x)(3/2 x 1/2x)=(2 x)(4/2 x) Distribute it over, and: 3 2 x/2x = 4

OpenStudy (anonymous):

Ok so is my part #1 answer correct? And is part #3 answer correct except for solving it? And for part #2 I changed it to using the LCD method, but can you tell me why we would use the LCD method?

OpenStudy (mathstudent55):